annette invested 8,400 in bank A at a certain interest rate per year and 7000 in bank B at an interest rate 1.5% higher. If the total income from the two investments is 751.80 at what interest rate was each amount invested?
Answer by solve_for_x(190)

Let x represent the rate at Bank A.

The rate at Bank B is then x + 1.5.

The total interest earned is then:

(8400)(x/100) + (7000)(x + 1.5)/100 = 751.80

Multiplying through by 100 gives:

8400x + 7000x + 10500 = 75180

15400x + 10500 = 75180

15400x = 75180 - 10500

15400x = 64680

x = 64680 / 15400

x = 4.2

Solution: The rate at Bank A is 4.2%, and the rate at Bank B is 4.2% + 1.5% = 5.7%


Check:

8400(0.042) + 7000(0.057) = 352.80 + 399.00 = 751.80